OpenStudy (erinweeks):

Find the solutions of the equation. 1/2x2 - x + 5 = 0

OpenStudy (erinweeks):

A. \[1 \pm \sqrt{9}i \] B. \[1 \pm \sqrt{11}i\] C. \[-1 \pm \sqrt{9}i\] D. \[-1 \pm \sqrt{11}i\]

OpenStudy (anonymous):

it can be rewritten as:: \[1-2x ^{3}+10x ^{2}=0\] or \[2x ^{3}-10x ^{2}-1=0\]

OpenStudy (erinweeks):

now could you show me the steps to solve that?

OpenStudy (anonymous):

find out the first root by hit and trial method

OpenStudy (erinweeks):

im not really sure how to do that

jimthompson5910 (jim_thompson5910):

1/2x^2 - x + 5 = 0 x^2 - 2x + 10 = 0 ... multiply every term by 2 Now use the quadratic formula to solve for x \[\Large x = \frac{-b\pm\sqrt{b^2-4ac}}{2a}\] \[\Large x = \frac{-(-2)\pm\sqrt{(-2)^2-4(1)(10)}}{2(1)}\] I'll let you finish. Tell me what you get

OpenStudy (erinweeks):

i got b some how.

jimthompson5910 (jim_thompson5910):

b is incorrect, try it again and post your work if you can

OpenStudy (erinweeks):

im not sure how to really solve it but im trying. i simplified the numbers under the radicand and got 33? then did 2(1) = 2 IM SO CONFUSED -_-

jimthompson5910 (jim_thompson5910):

\[\Large x = \frac{-b\pm\sqrt{b^2-4ac}}{2a}\] \[\Large x = \frac{-(-2)\pm\sqrt{(-2)^2-4(1)(10)}}{2(1)}\] \[\Large x = \frac{2\pm\sqrt{4-(40)}}{2}\] \[\Large x = \frac{2\pm\sqrt{-36}}{2}\] \[\Large x = \frac{2+\sqrt{-36}}{2} \ \text{or} \ x = \frac{2-\sqrt{-36}}{2}\] \[\Large x = \frac{2+6*i}{2} \ \text{or} \ x = \frac{2-6*i}{2}\] \[\Large x = 1+3*i \ \text{or} \ x = 1 - 3i\]

jimthompson5910 (jim_thompson5910):

now because sqrt(9) = 3 we can write the answer as \[\Large x = 1\pm \sqrt{9}*i\]

OpenStudy (erinweeks):

Omg Jim thank you so much.. I understand now (:

jimthompson5910 (jim_thompson5910):

you're welcome